OpenStudy (anonymous):

Can someone Please help me. And this is going to take a while so whoever helps me I want and need them to have time because I don't want either one of us rushing.. It's just a lot of algebra problems.

OpenStudy (anonymous):

ok what are the problems?

OpenStudy (anonymous):

okay how many do you want me to give at a time?

OpenStudy (anonymous):

all of them

OpenStudy (anonymous):

if u have them on a document send me the document or whatever file format u prefer

OpenStudy (anonymous):

okay hold on. It'll take about 3 minutes.

OpenStudy (anonymous):

ok not to worry i am solving another problem

OpenStudy (anonymous):

ok I have a question. like the first problem on my worksheet is a fraction and the only way you can see the fraction is if I draw it. So do you think I should just send it to you regular? And then when I have like fractions I can draw them for you.

OpenStudy (anonymous):

This is the first one. What is the sum? \[\frac{ 7 }{ 2 }+ \left(\begin{matrix}-3 \\ 5\end{matrix}\right)\]

OpenStudy (anonymous):

whichever way u prefer suit urself i will solve them all

OpenStudy (anonymous):

you mean

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } \frac{ 29 }{ 10 } \frac{ -21 }{ 10 } \frac{ 41 }{ 10 }\]

OpenStudy (anonymous):

\[\frac{ 7 }{ 2 }+(\frac{ -3 }{ 5 })\]

OpenStudy (anonymous):

Yeah I meant that. Sorry

OpenStudy (anonymous):

\[\frac{ 35-6 }{ 10 }=\frac{ 29 }{ 10 }=2.9\]

OpenStudy (anonymous):

Oh okay. I see. here's my second one. Which property is illustrated by the statement? (4 + 6.1) + 7 = 4 + (6.1 + 7) Associative Property of Addition Inverse Property of Multiplication Commutative Property of Addition Inverse Property of Addition

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } + \frac{ 29 }{ 10 } +\frac{ -21 }{ 10 } + \frac{ 41 }{ 10 }\]

OpenStudy (anonymous):

is that what you mean

OpenStudy (anonymous):

why don't u take a picture of that document with ur phone and send me the whole problem sets

OpenStudy (anonymous):

Okay I have a i phone but if I take a picture of it with my phone how am I gonna get it to you?

OpenStudy (anonymous):

well save the picture transfer the picture to ur pc using a cable

OpenStudy (anonymous):

okay hold on :)

OpenStudy (anonymous):

im gonna try it this way. 2. Which property is illustrated by the statement? (4 + 6.1) + 7 = 4 + (6.1 + 7) (1 point)Associative Property of Addition Inverse Property of Multiplication Commutative Property of Addition Inverse Property of Addition 3. To which subsets of the real numbers does the number belong? (1 point)rational numbers irrational numbers whole numbers, integers, rational numbers whole numbers, natural numbers, integers 4. What is the algebraic expression for the following word phrase: the product of 5 more than p and 7? (1 point)5 + p + 7 7(p + 5) 7 • p • 5 5. What is the simplified form of the expression ? (1 point)121 11 6. Evaluate , for x = and y = (1 point) 7. Simplify the expression . (1 point) 8. What is the simplified form of the expression 3[14 ÷ (42 – 32) – 5]? (1 point)–9 27 21 6 9. Is (5, 17) a solution of the equation y = 4x + 3? (1 point)yes no 10. Solve the equation. 3m – 4 = 26 (1 point)m = 30 m = m = 10 m = 17 11. Solve the equation. –15 = (1 point)a = –3 a = –20 a = –10 a = –75 12. Solve the equation. 3x – 10 = 3(x – 4) + 2 (1 point)x = 8 x = 0 x = infinitely many solutions x = no solution 13. Amy and Mike are going to an amusement park. They each bought ice cream, and Mike got a soda. Amy had a $10 coupon to put toward the cost, and Mike paid the rest, which came to $30.25. Admission for one person costs $18.00, and ice cream costs $1.50. How much does a soda cost at the amusement park? (1 point)$1.25 $1.50 $0.75 $2.10 14. Solve the equation. 7(2b – 5) + 3 = 10 (1 point)b = 42 b = b = b = 3 15. Suppose U = {1, 2, 3, 4, 5, 6, 7, 8} is the universal set and P = {1, 2, 3, 4}. What is P'? (1 point){5, 6, 7, 8} {1, 2, 3, 4, 5, 6, 7, 8} {1, 2, 3, 4} cannot be determined 16. Solve. 8x – 5 ≥ 27 (1 point)x ≥ 4 x ≥ x ≤ 4 x ≤ 17. Your goal is to save at least $360.00 over the next 6 weeks. How much money must you save each week in order to meet that goal? Write and solve an inequality. (1 point)6 + x ≥ 360; x ≥ 354 60x ≤ 360; x ≤ 10 ≤ 360; x ≤ 2,160 6x ≥ 360; x ≥ 60 18. A student scored 78 and 95 on his first two quizzes. Use a compound inequality to find the possible values for a third quiz score that would give him an average between 80 and 90, inclusive. (1 point)67 ≤ x ≤ 97 78 ≤ x ≤ 95 80 ≤ x ≤ 90 86 ≤ x ≤ 87 19. Of 256 consumers polled, some like only thriller movies, some prefer just comedies, and some like both. If 78 people like only thrillers and 116 like both, how many people prefer only comedies? (1 point)62 38 140 194 20. Choose the correct solution graph for the inequality. –3x > 24 (1 point) 21. Choose the correct solution graph for the inequality. 12x + 4 ≥ 16 or 3x – 5 ≤ –14 (1 point) 22. Solve. (1 point)x = – and x = x = and x = – x = – and x = – x = and x = 23. A new book is published and the table shows the number of people who showed up for a book signing in the first, second, third, and fourth month of its release. Which graph could represent the data shown in the table? Month Number of People 1 64 2 134 3 216 4 343 (1 point) 24. The table shows the relationship between two variables. Is the relationship a linear function? x y 1 1 2 –3 3 –7 4 –11 (1 point)yes no 25. The ordered pairs (1, 5), (2, 25), (3, 125), (4, 625), and (5, 3125) represent a function. What is a rule that represents this function? (1 point)y = x5 y = 5x y = 5x y = x + 5 26. Suppose noodle soup is on sale for $0.75 a can and you have a coupon for $0.50 off your total purchase. Write a function rule for the cost of n cans of soup. (1 point)C(n) = 0.75n – 0.5 C(n) = 0.5n – 0.75 C(n) = (0.75 + n) – 0.5 C(n) = (0.5 + n) – 0.75 Note: This item has been reviewed and is scheduled to be updated. All students will receive full credit for any response to the following. 27. Determine if the mapping diagram that represents a relation is a function. (1 point)The relation is a function. The relation is not a function. 28. Tell whether the sequence is arithmetic. If it is, what is the common difference? –12, –7, –2, 3, . . . (1 point)yes; 5 yes; 8 yes; 3 no 29. What is the graph of the function rule? y= |2 x | + 1 (1 point) 30. For the data in the table, does y vary directly with x? If it does, write an equation for the direct variation. x y 3 6 6 18 8 24 (1 point)yes; y = 2x yes; y = 3x yes; y = 4x no; y does not vary directly with x 31. Match the equation with its graph. –x + y = (1 point) 32. Find the slope of the line. (1 point) – –3 5 33. What is the slope of the line that passes through the pair of points? (–2, 7), (18, 1) (1 point) – – – 34. What is the slope of the line? (1 point)–2 0 undefined 2 35. Write an equation in point-slope form for the line through the given point with the given slope. (–7, 9); m = (1 point)y + 7 = (x – 9) y – 7 = (x + 9) y – 9 = (x + 7) y + 9 = (x – 7) 36. The table shows the height of a tree as it grows. What equation in slope-intercept form gives the tree's height at any time? Time (months) Height (inches) 2 14 4 23 6 32 8 41 (1 point)y = 9x + 2 y = x + 5 y = x y = 5x + 37. Write y = –x – 3 in standard form using integers. (1 point)2x + 5y = –15 2x + y = –3 2x + 5y = –3 x + y = –15 38. Tell whether the lines for the pair of equations are parallel, perpendicular, or neither. y = –x + 1 2x – 3y = –3 (1 point)parallel perpendicular neither 39. Write the equation of a line that is perpendicular to the given line and that passes through the given point. y + 2 = (x – 5); (–4, 3) (1 point)y = –3x + 15 y = –3x – 9 y = –4x + 3 y = –x – 11 40. Graph y = │x + 2│ (1 point) 41. Which graph represents the solution to the given system? y = –2x + 1 y = x + 6 (1 point) 42. How many solutions does the system have? 3x + 12y = 20 y = –x + (1 point)one two infinite none 43. Which inequality represents the graph below? (1 point)y ≥ 5 x ≥ 5 y ≤ 5 x ≤ 5 44. Bradley and Ryan are having a contest to see who can collect the most bumper stickers. Bradley started with 23 bumper stickers and plans to add 4 each month. Ryan has 9 bumper stickers and plans to add 6 each month. After how many months will they have the same number of bumper stickers? (1 point)10 months 2 months 7 months 51 months 45. Which system is represented by the graph? (1 point) 8x – y = 1 8x – y = –3 x – 8y = 1 x – 8y = –3 8x – y = –1 8x – y = 3 x – 8y = –1 x – 8y = 3 46. What is the solution of the system? y = 14x – 1 y = –11x + 4 (1 point)(5, –51) 47. Which graph represents the given inequality? y > 4x – 2 (1 point) 48. What system of inequalities is represented by the graph? (1 point) y ≤ 3x – 4 5x + y ≤ 2 y ≥ 3x – 4 5x + y ≤ 2 y ≥ 3x – 4 5x + y ≥ 2 y ≤ 3x – 4 5x + y ≥ 2

OpenStudy (anonymous):

ok simple but u should know each one of those have more than one method to solve them ok

OpenStudy (anonymous):

ok. and you know how some of these are missing numbers? well thats because there fractions. and I guess wen we get to a problem that I need to send you a picture for it I will

OpenStudy (anonymous):

ok not to worry about that

OpenStudy (anonymous):

okay good

OpenStudy (anonymous):

by the way some of those question are already answered

OpenStudy (anonymous):

questions*

OpenStudy (anonymous):

oh they are. I don't know how to tell or not :(

OpenStudy (anonymous):

Do you want to start working on it now?

OpenStudy (anonymous):

hehehe i will highlight them with red color

OpenStudy (anonymous):

have a question.I know we don't know each other at all. But do you have a i phone? cause if you have a i phone I could send you the pictures off of my phone and that would make it soooo easy!!!

OpenStudy (anonymous):

no don't use any smart phones

OpenStudy (anonymous):

plus i am from the UK not in the state

OpenStudy (anonymous):

those smart phones dulls minds

OpenStudy (anonymous):

its a waste of time

OpenStudy (anonymous):

they always break!! Mine is broke is some places now :( but anyway okay lets work on this because my teacher is giving me all today to finsish this and if i dont finish it then im gonna fail my first semester of algebra :(((((

OpenStudy (anonymous):

no u won't i will finish the job and explain it for u

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

if u look at the what is written after (1 point) u'll see the answers they already exists

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

2. Which property is illustrated by the statement? (4 + 6.1) + 7 = 4 + (6.1 + 7) (1 point) Associative Property of Addition

OpenStudy (anonymous):

(a+b)+c = a +(b+c) this called Associative Property of addition so never forget that, its important for ur fundamentals

OpenStudy (anonymous):

okay I thought thats what it was. just wasntt sure but thanks for telling me why.

OpenStudy (anonymous):

ok as for Q3 the number is missing

OpenStudy (anonymous):

To which subsets of the real numbers does the number\[\frac{ 1 }{ 3 }\] belong?

OpenStudy (anonymous):

rational numbers irrational numbers whole numbers, integers, rational numbers whole numbers, natural numbers, integers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/3 belongs to the subset of rational numbers

OpenStudy (anonymous):

so fractions belongs to rational numbers subset but sqrt (2) belongs to irrational numbers

OpenStudy (anonymous):

ok this is 4. What is the algebraic expression for the following word phrase: the product of 5 more than p and 7? 5 + p + 7 7(p + 5) 7x p x 5 \[\frac{ p + 5 }{ 7 }\]

OpenStudy (anonymous):

there is a missing part for question 4

OpenStudy (anonymous):

I just drew it right. :)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

that question is a wrong question because he did not define p

OpenStudy (anonymous):

we will live it for now till i fix it let's move on

OpenStudy (anonymous):

as for Q5 121 the simplified term is 11

OpenStudy (anonymous):

Q6 is missing

OpenStudy (anonymous):

are u there?

OpenStudy (anonymous):

yes hold on for about a minute

OpenStudy (anonymous):

okay im gonna write down question 6 for you.

OpenStudy (anonymous):

just give me the missing numbers

OpenStudy (anonymous):